From the measured values shown, determine if the measurements are within the required precision. Show all computations. Allowable precision is 1/5000.Determine the most probable value of the interior angles.If the measured distance are 220.87,220.90,220.93,220.81 and 220.89.


1
Expert's answer
2021-01-21T06:35:11-0500

Σ distance = 220.87 + 220.90 + 220.93 + 220.81 + 220.89

=1,104.4

∴ Weighted arithmetic mean ="\tfrac{1104.4}{5}"

= 220.88


the most probable value = 220.88 +"\frac{1}{5000}" or 220.88 -"\frac{1}{5000}"

=220.8802 or 220.8798